from a group of 6 men and 4 women, a committee of 3 is to be selected at random. find P(at least 2 women)
Found 2 solutions by jim_thompson5910, edjones:
Answer by jim_thompson5910(35256)

Simply find the probability of 2 women or 3 women. In other words,


P(at least 2 women) = P(2 women) + P(3 women)

Answer by edjones(8007)

The probability of selecting a man (m) is .6 and a woman is .4
the probability of all women is w^3=.4^3=.064
2/3 being women is 3mw^2=3*.6*.4^2=.288 [the third term of (m+f)^3].
.288+.064=.352
.
Ed